OpenStudy (anonymous):

Solve: 5(x + 8) = 7(x − 2)

OpenStudy (anonymous):

x = 27

OpenStudy (anonymous):

2x=54 x=27

OpenStudy (anonymous):

5x +40 = 7x -14 5x-7x=-14 - 40 -2x = -54 x = 27

OpenStudy (anonymous):

1. Solve: x − 5 = 12 2. Solve: −3x − 5 = 16

OpenStudy (anonymous):

(1) x=12+5 x= 17

OpenStudy (anonymous):

(2) -3x=16 +5 -3x= 21 x=-7

OpenStudy (anonymous):

Show them how to solve it for themselves, don't just give them then answers.